a radio active element is placed in a air free zone what will be change of decay rate per second of the element

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

decay rate=dy/dt=-λt. change in decay rate per second= rate of change of decay rate=(d/dt)(dy/dt) =(d/dt)(-λt) =-λ second-1